Two garden hoses of different diameter are needed to reach a flower bed. Water flows with a speed of 1.2m/s through the bottom section of a garden hose that goes up a step 20.0 cm high. The cross-sectional area of the hose at the top of the step is half that at the bottom of the step. If the water pressure is 143 kPa at the bottom of the step, what is its pressure and velocity at the top of the step?
Answer #1

Volume flow velocity (A*v) of water will be same everywhere.

pi x r^2 x 1.2 = pi x (r/2)^2 x v

v = 4 x 1.2 =4.8 m/s ...........Ans

Using bernoulli's equation,

P + row*gh + row*v^2 /2 = constant

for bottom = for top

143 x 10^3 + 1000 x 9.81 x 0 + 1000 x 1.2^2 /2 = P + 1000 x 9.81 x 0.20 + 1000x 4.8^2 /2

P = 130238 Pa .......Ans